# Implementing Errors on Errors: Bayesian vs Frequentist

### v1

@@ 4 Comparison between D’Agostini and Cowan
[…]
@@
We see that the choice ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}$, as suggested in Eq. (25), provides a better match with the result from Cowan’s frequentist approach.888 Cowan also touches on related Bayesian approaches in his talk [7], noting that choosing ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}$ yields a gamma distribution with unit expectation and adjustable variance. His reference to the inverse gamma distribution, in comparison with D’Agostini’s approach, suggests a perspective not fully aligned with the formulation considered here, which relies exclusively on the gamma distribution. The corresponding PDFs are shown in Fig. 1 with the values ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.2$, $0.5$, and $1$, namely, ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=6.25$, $1$, and $0.25$, respectively. Accordingly, the corresponding numbers of virtual repetitions are ${\color[rgb]{0.03,1,0.25}n}=13.5$, $3$, and $1.5$, respectively. Among these, the choice ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.5$ corresponds to our version of D’Agostini’s criterion $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]$, leading to Eq. (27).
@@ 5 Summary and discussion
Before closing, we note possible directions for future work. The hierarchical Bayesian model with hyperpriors on variance parameters [8] has been proposed as a way to address issues like “sub-sampling” [9], where iterative rescaling can bias the combined central value. It would be worthwhile to examine whether this issue can also be resolved within a frequentist approach, and whether the correspondence identified in this Letter continues to hold in that context. In parallel, it would also be valuable to explore how the frameworks discussed here perform when applied to real-world data sets [10], such as those underlying the $W$ boson mass [11], the muon $g-2$ anomaly [12], the neutron lifetime [13] (and references therein), and the $\sigma_{8}$ tension in cosmology [14]. Such applications may offer not only practical validation but also a clearer understanding of the statistical structure behind persistent discrepancies in precision measurements.

### v2

@@ 4 Comparison between D’Agostini and Cowan
$\underset{\text{D'Agostini's Bayesian}}{{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}={1\over{\color[rgb]{1,0.2,0}\hat{r}_{i}^{2}}}={{\color[rgb]{0,0.5,1}s_{i}^{2}}\over{\color[rgb]{1,0.2,0}\hat{\sigma}_{i}^{2}}}}$ $\stackrel{{\scriptstyle\text{correspond}}}{{\longleftrightarrow}}\underset{\text{Cowan's frequentist}}{{\color[rgb]{0,0.5,1}\hat{w}_{i}}\over{\color[rgb]{1,0.2,0}v_{i}}}.$ (44)
@@
We see that the choice ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}$, as suggested in Eq. (25), provides a better match with the result from Cowan’s frequentist approach.101010 Cowan also touches on related Bayesian approaches in his talk [8], noting that choosing ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}$ yields a gamma distribution with unit expectation and adjustable variance. His reference to the inverse gamma distribution, in comparison with D’Agostini’s approach, suggests a perspective not fully aligned with the formulation considered here, which relies exclusively on the gamma distribution. The corresponding PDFs are shown in Fig. 1 with the values ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.2$, $0.5$, and $1$, namely, ${\color[rgb]{0.03,1,0.25}k}={\color[rgb]{0.03,1,0.25}\lambda}=6.25$, $1$, and $0.25$, respectively. Accordingly, the corresponding numbers of virtual repetitions are ${\color[rgb]{0.03,1,0.25}n}=13.5$, $3$, and $1.5$, respectively. Among these, the choice ${\color[rgb]{0.03,1,0.25}\varepsilon}=0.5$ corresponds to our version of D’Agostini’s criterion $\operatorname{V}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]=\operatorname{E}^{2}\mathchoice{\!}{\!}{}{}\left[{\color[rgb]{1,0.2,0}\hat{\omega}_{i}}\right]$, leading to Eq. (27).
@@ 5 Summary and discussion
Before closing, we note possible directions for future work. The hierarchical Bayesian model with hyperpriors on variance parameters [9] has been proposed as a way to address issues like “sub-sampling” [10], where iterative rescaling can bias the combined central value. It would be worthwhile to examine whether this issue can also be resolved within a frequentist approach, and whether the correspondence identified in this Letter continues to hold in that context. In parallel, it would also be valuable to explore how the frameworks discussed here perform when applied to real-world data sets [11], such as those underlying the $W$ boson mass [12], the muon g2g-2 anomaly [4],111111 Applications to the muon g2g-2 have already been considered in Ref. [5]; see also recent developments on the data-theory comparison in Refs. [13, 14]. the neutron lifetime [15] (and references therein), and the $\sigma_{8}$ tension in cosmology [16]. Such applications may offer not only practical validation but also a clearer understanding of the statistical structure behind persistent discrepancies in precision measurements.